Please help solve this f(x)=x^2+3x+2
Answer by nerdybill(7384)

f(x)=x^2+3x+2
to find the "roots", we set f(x) to zero and solve for x:
0=x^2+3x+2
factoring the right:
0 = (x+1)(x+2)
set each term on the right to zero and solve for x we get
x = {-1, -2}